# Identifying Spatiotemporal Patterns in Opioid Vulnerability: Investigating the Links Between Disability, Prescription Opioids and Opioid-Related Mortality

### v2

@@ 1   Introduction
For data estimation and prediction, the cornerstone of our analysis utilizes a Kalman filter [14, 15]. The Kalman filter uses a series of observations to estimate unknown parameters and has been applied in many diverse settings, from navigating astronauts to the moon to real-time vehicle tracking [16, 17]. Such applications most commonly occur in temporal statistical modeling, as the method traditionally lacks a spatial framework. To overcome this limitation, we propose a novel approach to augmenting the Kalman filter with a spatial component. This yields a principled framework to capture the complex geographical interrelations among counties within the United State’s diverse landscape, thereby improving the accuracy of our opioid-related data estimations and predictions.
To define opioid vulnerability profiles, we create heat maps of our filter’s predicted rates across the nation’s counties and identify the hotspots. In this context, hotspots are defined on a year-by-year basis as counties with rates in the top 5% nationally. Multiple contiguous hotspot counties are referred to as clusters of hotspots and indicate the most vulnerable areas in the nation.
@@ 1.1   Related works
The intertwined dynamics of opioid mortality, prescription rates, and social vulnerability has been explored in previous research [21, 22, 23, 24]. Such studies have concluded that the crisis’s roots extend beyond prescription practices to broader socioenvironmental issues and advocate for addressing structural addiction determinants [25, 26]. Specifically, factors such as disability, minority status, lack of health insurance, and low income have been identified as significant predictors of opioid misuse and mortality [27, 28, 29]. Building upon such previous insights, our study offers novel geospatial evidence to the discussion.
@@ 2.1   Data
This study examines the rates of opioid-related mortality, opioid prescription dispensing, and disability across the United States from 2014 to 2020. It is important to note that each dataset was collected differently: the opioid-related mortality rates are measured per 100,000 persons, the opioid prescription dispensing rates are measured per 100 persons, and the disability rank ordered rates are measured as percentile ranks from 0 to 100. Throughout our study period, the county structure of the United States changed due to the formation of new counties and the reconfiguration of existing ones [42]. However, the Kalman filter requires a fixed state space for its predictions, so our data has been curated to reflect the 2020 county structure, which includes 3,143 counties. This step ensures that we provide the most up-to-date and cohesive representation of the national county landscape with respect to the data.
@@ 2.2   Kalman filter model
The Kalman filter operates on two principal equations: the state update equation, which models the evolution of a system’s state over time, and the observation equation, which links the system’s true state to observed data. These enable the Kalman filter to refine its estimates by updating previous predictions with new data. Opioid-related county-level data inherently possess both geospatial and temporal dimensions, making our data well-suited for the Kalman filter. Its inherent temporal nature allows it to leverage the evolution of opioid-related data over time, uncovering temporal patterns and trends. The spatial dimension is incorporated through the filter’s covariance matrix, capturing the spatial correlations between counties in the United States. A detailed mathematical exposition of the Kalman filter algorithm is provided in the appendix.
@@
The normal random vector $\eta_{t}$ represents the measurement error in data collection, reporting, or other discrepancies from the true state. $R$ is the observation covariance matrix and quantifies the uncertainty in the data measurements. To choose a data uncertainty level for our study, we conducted a sensitivity analysis. This analysis examined various uncertainty levels, specifically 1%, 3%, and 5%, by evaluating the respective changes in the model’s predictive performance. Optimal predictive accuracy was found with a 1% data uncertainty level. Consequently, a 1% uncertainty assumption was adopted for all data measurements throughout our study. The results of this sensitivity analysis are summarized in the appendix.
@@ 3.1   Efficacy analysis of the filter’s performance in the 2020 prediction year
Presented here are the results for our spatial Kalman filter which was initialized with 2014 data and trained on data from 2015 to 2019 before generating predictions for 2020. Training, in this context, means allowing the filter to refine its estimates using the observed data; while its estimates were made without further refinement in the 2020 prediction year. We chose to predict for only a single year to prevent the accumulation of large errors inherent in multi-year forecasting. Long-term forecasting is accommodated by the filter’s design however, and investigated in the appendix. In this investigation, the efficacy of multi-year predictions from the filter is examined by comparing its performance when trained on progressively less data to that of the fully trained filter discussed here. Presented in this section are the efficacy results for the filter’s 2020 predictions. For the remaining years in the study, the efficacy metrics are summarized in the appendix. In addition, the corresponding accuracy maps and error histograms can be found in the supplementary materials.
@@
For opioid-related mortality rates, measured per 100,000 persons, the filter achieved an average general accuracy of 93.62%. The national accuracy distribution is displayed in Figure 1. This map is characterized predominantly by green hues, indicating high predictive accuracy across the nation. A distinct region of lower accuracy, marked by yellow and red hues, is found in the southern part of West Virginia. This region experienced significant jumps in mortality rates from 2019 to 2020. These large jumps were difficult for the filter to predict and led to these distinguishable errors. The range of errors for this dataset, shown in fig. 2, illustrates that our model’s mortality rate estimates can deviate from the actual rates by an average of 5.57 deaths per 100,000 persons, with a maximum deviation of 87.27 deaths per 100,000 persons in the worst case. Additionally, for this dataset, the filter achieved a hotspot accuracy of 71.50%. The hotspots which were correctly identified and those that were missed are shown in fig. 3.
Figure 1: General accuracy maps for the 2020 (a) mortality, (b) disability, and (c) dispensing rate predictions. The maps are color-coded to represent increasing intervals of accuracy by 5%, starting from dark red for the least accurate predictions, and progressing to dark green for the most accurate predictions.
@@
Figure 3: Hotspot accuracy maps for the 2020 (a) mortality, (b) disability, and (c) dispensing rate predictions. Orange colored counties represent accurately predicted hotspots, whereas those in black mark hotspots missed by our predictions.
For the disability rank ordered rates, measured as percentile ranks from 0 to 100, our spatial Kalman filter demonstrated excellent predictive performance. The range of errors, shown in fig. 2, demonstrates that our model’s disability rank ordered rate estimates can deviate from the actual rank ordered rates by an average of 1.57 percentile rank points, with a maximum deviation of 15.85 percentile rank points in the worst case. This efficacy of the filter is also highlighted by its high hotspot accuracy of 95.35%. The correctly predicted hotspots and those that were missed are shown in fig. 3. On the other hand, the average general accuracy score of 90.08% may not fully convey the filter’s effectiveness on this dataset. This is because the maximum annual error of 15.85 used for normalization on this dataset results in even relatively small errors appearing more pronounced. This is visually evident in fig. 1; where we observe more yellow tones, denoting this magnification of smaller errors.
For the prescription dispensing rates, measured per 100 persons, the filter’s predictions attained an average general accuracy of 91.69%. The national accuracy distribution is visualized in fig. 1. This map is predominantly green; but upon close inspection, two red counties stand out: Union, South Dakota, and Montour, Pennsylvania. These counties experienced an extraordinary increase of over 350% in their dispensing rates from 2018 to 2019. Based on this surge, the filter predicted similarly high increases from 2019 to 2020. However, the increases in 2020 were far less dramatic, leading to these significant prediction errors. The range of errors for this dataset, shown in Figure 2, illustrates that our model’s dispensing rate estimates can deviate from the actual dispensing rates by an average of 20.81 prescriptions per 100 persons, with a maximum deviation of 250.27 prescriptions per 100 persons in the worst case. Additionally, the filter’s hotspot accuracy on this dataset was 68.05%. The correctly identified hotspots and those that were missed are shown in fig. 3.
@@ 3.2   National opioid vulnerability profiles
In this section, we present the vulnerability profiles uncovered by our spatial Kalman filter’s predictions. The heat maps and hotspot maps for the rates of mortality, disability and dispensing are showcased in fig. 4 through fig. 6, respectively. To illustrate the temporal evolution of each factor, the figures present the most significant maps. For both the mortality and disability rate predictions, we highlight maps from the years 2014, 2017, and 2020, while for the dispensing rate predictions, we focus on the years 2014, 2018, and 2019. Maps for the remaining years in the study can be found in the images folder of the supplementary materials.
Figure 4: Heat maps and hotspot maps for the spatial Kalman filter’s opioid-related mortality rate predictions in (a) 2014, (b) 2017, and (c) 2020. Counties with missing data are colored in black.
Figure 5: Heat maps and hotspot maps for the spatial Kalman filter’s disability rank ordered rate predictions in (a) 2014, (b) 2017, and (c) 2020. Counties with missing data are colored in black.
Figure 6: Heat maps and hotspot maps for the spatial Kalman filter’s prescription opioid dispensing rate predictions in (a) 2014, (b) 2018, and (c) 2019. Counties with missing data are colored in black.
@@ 4   Discussion
Leveraging the accurate predictions from our spatial Kalman filter, we analyzed the spatiotemporal patterns of the opioid crisis for the years 2014 through 2020 via constructed vulnerability profiles for three key factors: opioid-related mortality rates, opioid-dispensing rates, and disability rank ordered rates. The analysis of profiles for the mortality rates revealed widespread national vulnerability where the Appalachian region was repeatedly distinguished as the nation’s most vulnerable area to opioid-related mortality rates. Correspondingly, the analysis examining the disability rank ordered rate profiles also uncovered widespread vulnerability where Appalachia was again repeatedly pinpointed as the nation’s most vulnerable area to disability rank ordered rates. This alignment in the patterns for both disability and mortality profiles suggests that amidst the general fabric of the opioid crisis, the disabled subpopulation may be more at risk of opioid-related mortality than the general population. This finding corroborates previous work [27, 25, 50], and underscores the disproportionate outcomes of the opioid crisis.
@@ 5   Conclusion
In this study, we utilized the predictions from our spatial Kalman filter to investigate opioid vulnerability profiles and detect critical spatiotemporal patterns. The methodologies employed in this research could be further adapted and refined by future studies to investigate a broader range of phenomena incorporating both spatial and temporal dimensions. Moreover, future research could leverage the filter’s potential for long-term forecasting to predict outcomes over extended periods.
This study does have certain limitations to note, particularly the restricted range of years for which consistent data is available across all three datasets. The dispensing rates contained a notable amount of missing data, which posed challenges for data handling. Moreover, the final year of this study’s time frame coincides with the onset of the COVID-19 pandemic, confounding the pandemic’s broader societal effects with the societal drivers of opioid abuse. These effects may limit the comparability of 2020 data to prior years and affect the generalizability of findings from that year.

### v3

@@ 1   Introduction
For data estimation and prediction, the cornerstone of our analysis utilizes a Kalman filter [14, 15]. The Kalman filter uses a series of observations to estimate unknown parameters and has been applied in many diverse settings, from navigating astronauts to the moon to real-time vehicle tracking [16, 17]. Despite its versatility, such applications primarily focus on temporal statistical modeling, as the method traditionally lacks a spatial framework. To overcome this limitation, we propose a novel approach to augmenting the Kalman filter with a spatial component. This yields a principled framework to capture the complex geographical interrelations among counties within the United State’s diverse landscape, thereby improving the accuracy of our opioid-related data estimations and predictions.
To define opioid vulnerability profiles, we create heat maps of our filter’s predicted rates across the nation’s counties and identify the hotspots. In this context, hotspots are defined on a year-by-year basis as counties with rates in the top 5% nationally. Multiple contiguous hotspot counties are referred to as clusters of hotspots and indicate the most vulnerable areas in the nation. Although many studies have examined spatiotemporal variations in opioid-related outcomes, most have concentrated on specific states or localized areas, limiting their ability to capture national patterns and trends [18, 19, 20, 21, 22, 23]. Additionally, much of this research draws from clinical data sources, such as hospital records or autopsy reports, which, while valuable, may not fully capture the broader public health dynamics at play. To address these gaps, we examine the opioid crisis from a nationwide public health perspective, leveraging county-level data to thoroughly investigate community-level vulnerabilities across the United States.
@@ 1.1   Related works
The intertwined dynamics of opioid mortality, prescription rates, and social vulnerability has been explored in previous research [27, 28, 29, 30]. Such studies have concluded that the crisis’s roots extend beyond prescription practices to broader socioenvironmental issues and advocate for addressing structural addiction determinants [31, 32]. Specifically, factors such as disability, demographic status, lack of health insurance, and low income have been identified as significant predictors of opioid misuse and mortality [33, 34, 35]. Building upon such previous insights, our study offers novel geospatial evidence to the discussion.
@@ 2.1   Data
This study examines the rates of opioid-related mortality, opioid prescription dispensing, and disability across the United States from 2014 to 2020. It is important to note that each dataset was collected differently: the opioid-related mortality rates are measured per 100,000 persons, the opioid prescription dispensing rates are measured per 100 persons, and the disability rank ordered rates are measured as percentile ranks from 0 to 100. Throughout our study period, the county structure of the United States changed due to the formation of new counties and the reconfiguration of existing ones [47]. However, the Kalman filter requires a fixed state space for its predictions, so our data has been curated to reflect the 2020 county structure, which includes 3,143 counties. This step ensures that we provide the most up-to-date and cohesive representation of the national county landscape with respect to the data. Additionally, at most four counties were added or removed in any given year throughout the study period, meaning this curation to the 2020 structure had minimal impact on our overall results.
@@ 2.2   Kalman filter model
The Kalman filter operates on two principal equations: the state update equation, which models the evolution of a system’s state over time, and the observation equation, which links the system’s true state to observed data. These enable the Kalman filter to refine its estimates by updating previous predictions with new data. Opioid-related county-level data inherently possess both geospatial and temporal dimensions, making our data well-suited for the Kalman filter. Its inherent temporal nature allows it to leverage the evolution of opioid-related data over time, uncovering temporal patterns and trends. The spatial dimension is incorporated through the filter’s covariance matrix, capturing the spatial correlations between counties in the United States. A detailed mathematical exposition of the Kalman filter algorithm is provided in appendix A.
@@
We selected the exponential covariance function over alternatives, such as the Gaussian, because of its ability to better capture the complex spatial patterns observed in our data. Opioid-related outcomes often exhibit abrupt changes between neighboring counties that a Gaussian function, which enforces smoother spatial transitions, may fail to capture. The exponential decay function accommodates these sharp transitions while preserving long-range correlations, offering a better fit for the non-smooth spatial variations seen in the heat maps. Additionally, the Gaussian function’s light tails make it less effective at capturing rare events, which are important when modeling abrupt regional disparities in opioid-related outcomes.
In constructing the covariance matrix, we used distances between county centroids rather than adjacency-based connections to better capture continuous spatial relationships across U.S. counties. Adjacency-based models define spatial correlation strictly through shared borders, imposing a binary structure where counties are either connected or not, potentially oversimplifying complex spatial dependencies. This can be particularly limiting for opioid-related outcomes, which may be influenced by broader regional dynamics extending beyond immediate neighbors. By leveraging centroid distances between county population centers, our approach models gradual changes in spatial correlation across varying distances. Moreover, given the significant heterogeneity in county sizes and shapes across the United States, centroid-based distances provide a consistent and scalable measure of proximity.
@@
The normal random vector $\eta_{t}$ represents the measurement error in data collection, reporting, or other discrepancies from the true state. $R$ is the observation covariance matrix and quantifies the uncertainty in the data measurements. To choose a data uncertainty level for our study, we conducted a sensitivity analysis. This analysis examined various uncertainty levels, specifically 1%, 3%, and 5%, by evaluating the respective changes in the model’s predictive performance. Optimal predictive accuracy was found with a 1% data uncertainty level. Consequently, a 1% uncertainty assumption was adopted for all data measurements throughout our study. The results of this sensitivity analysis are summarized in appendix B.
@@ 3.1   Efficacy analysis of the filter’s performance in the 2020 prediction year
Presented here are the results for our spatial Kalman filter which was initialized with 2014 data and trained on data from 2015 to 2019 before generating predictions for 2020. Training, in this context, means allowing the filter to refine its estimates using the observed data; while its estimates were made without further refinement in the 2020 prediction year. We chose to predict for only a single year to prevent the accumulation of large errors inherent in multi-year forecasting. Long-term forecasting is accommodated by the filter’s design however, and investigated in appendix C. In this investigation, the efficacy of multi-year predictions from the filter is examined by comparing its performance when trained on progressively less data to that of the fully trained filter discussed here. In appendix D, we explore how shifting the initialization year affects the model’s predictions, assessing whether more recent training data improves forecasting accuracy. Presented in this section are the efficacy results for the filter’s 2020 predictions. For the remaining years in the study, the efficacy metrics are summarized in appendix E. In addition, the corresponding accuracy maps and error histograms can be found in the supplementary materials.
@@
For opioid-related mortality rates, measured per 100,000 persons, the filter achieved an average general accuracy of 94.00%. The national accuracy distribution is displayed in Figure 1. This map is characterized predominantly by green hues, indicating high predictive accuracy across the nation. A distinct region of lower accuracy, marked by yellow and red hues, is found in the southern part of West Virginia. This region experienced significant jumps in mortality rates from 2019 to 2020. These large jumps were difficult for the filter to predict and led to these distinguishable errors. The range of errors for this dataset, shown in fig. 2, illustrates that our model’s mortality rate estimates can deviate from the actual rates by an average of 5.20 deaths per 100,000 persons, with a maximum deviation of 86.61 deaths per 100,000 persons in the worst case. Additionally, for this dataset, the filter achieved a hotspot accuracy of 70.00%. The hotspots which were correctly identified and those that were missed are shown in fig. 3.
Figure 1: General accuracy maps for the 2020 (a) mortality, (b) disability, and (c) dispensing rate predictions. The maps are color-coded to represent increasing intervals of accuracy by 5%, starting from dark red for the least accurate predictions, and progressing to dark green for the most accurate predictions. Figures are best viewed online in color.
@@
Figure 3: Hotspot accuracy maps for the 2020 (a) mortality, (b) disability, and (c) dispensing rate predictions. Orange colored counties represent accurately predicted hotspots, whereas those in black mark hotspots missed by our predictions. Figures are best viewed online in color.
For the disability rank ordered rates, measured as percentile ranks from 0 to 100, our spatial Kalman filter demonstrated excellent predictive performance. The range of errors, shown in fig. 2, demonstrates that our model’s disability rank ordered rate estimates can deviate from the actual rank ordered rates by an average of 1.84 percentile rank points, with a maximum deviation of 15.97 percentile rank points in the worst case. This efficacy of the filter is also highlighted by its high hotspot accuracy of 94.19%. The correctly predicted hotspots and those that were missed are shown in fig. 3. On the other hand, the average general accuracy score of 88.49% may not fully convey the filter’s effectiveness on this dataset. This is because the maximum annual error of 15.97 used for normalization on this dataset results in even relatively small errors appearing more pronounced. This is visually evident in fig. 1; where we observe more yellow tones, denoting this magnification of smaller errors.
For the prescription dispensing rates, measured per 100 persons, the filter’s predictions attained an average general accuracy of 91.61%. The national accuracy distribution is visualized in fig. 1. This map is predominantly green; but upon close inspection, two red counties stand out: Union, South Dakota, and Montour, Pennsylvania. These counties experienced an extraordinary increase of over 350% in their dispensing rates from 2018 to 2019. Based on this surge, the filter predicted similarly high increases from 2019 to 2020. However, the increases in 2020 were far less dramatic, leading to these significant prediction errors. The range of errors for this dataset, shown in Figure 2, illustrates that our model’s dispensing rate estimates can deviate from the actual dispensing rates by an average of 21.02 prescriptions per 100 persons, with a maximum deviation of 250.67 prescriptions per 100 persons in the worst case. Additionally, the filter’s hotspot accuracy on this dataset was 68.64%. The correctly identified hotspots and those that were missed are shown in fig. 3.
For context on the model’s performance across the different datasets, table 3 in appendix F presents the summary statistics for each factor across the entire study period. These statistics highlight the differences in data variability within each dataset, which can help explain the observed differences in model efficacy across the different datasets. The disability rates, being percentile rank ordered, form a more uniform and less variable dataset. In contrast, the prescription dispensing rates are inherently more chaotic. As shown in table 3, the dispensing rates consistently exhibit the highest values and standard deviations each year. The variability within the mortality rates dataset falls somewhere between these two extremes, it displays more structure than the chaotic dispensing rates but lacks the ordered nature of the rank-ordered disability rates. These inherent differences in data distribution and variability likely contribute to the observed variation in model performance across the three factors.
@@ 3.2   National opioid vulnerability profiles
In this section, we present the vulnerability profiles uncovered by our spatial Kalman filter’s predictions. Before proceeding to the results, it is important to clarify how the predictions used to generate these vulnerability profiles were produced. For the training years (2014–2019), the spatial Kalman filter updates its forecasts by incorporating the actual data from each year, resulting in calibrated estimates. In contrast, the 2020 predictions represent raw forecasts, as the model generates them solely based on patterns learned from the 2015–2019 training period, without access to 2020 data for adjustment. In all cases, yearly data were fed to the model sequentially.
The heat maps and hotspot maps for the rates of mortality, disability and dispensing are showcased in fig. 4 through fig. 6, respectively. To illustrate the temporal evolution of each factor, the figures present the most significant maps. For both the mortality and disability rate predictions, we highlight maps from the years 2014, 2017, and 2020, while for the dispensing rate predictions, we focus on the years 2014, 2018, and 2019. Maps for the remaining years in the study can be found in the images folder of the supplementary materials.
Figure 4: Heat maps and hotspot maps depicting the spatial Kalman filter’s predicted opioid-related mortality rates for (a) 2014, (b) 2017, and (c) 2020. These maps represent model-generated estimates rather than observed data. Counties with missing data are colored in black. Figures are best viewed online in color.
Figure 5: Heat maps and hotspot maps depicting the spatial Kalman filter’s predicted disability rank-ordered rates for (a) 2014, (b) 2017, and (c) 2020. These maps represent model-generated estimates rather than observed data. Counties with missing data are colored in black. Figures are best viewed online in color.
Figure 6: Heat maps and hotspot maps depicting the spatial Kalman filter’s predicted prescription opioid dispensing rates for (a) 2014, (b) 2018, and (c) 2019. These maps represent model-generated estimates rather than observed data. Counties with missing data are colored in black. Figures are best viewed online in color.
@@ 4   Discussion
Leveraging the accurate predictions from our spatial Kalman filter, we analyzed the spatiotemporal patterns of the opioid crisis for the years 2014 through 2020 via constructed vulnerability profiles for three key factors: opioid-related mortality rates, opioid-dispensing rates, and disability rank ordered rates. Before analyzing these results in detail, it is important to validate our modeling approach by comparing the spatial Kalman filter to its generic counterpart. To ensure a fair evaluation, we assessed absolute errors rather than accuracy-based metrics. Accuracy measures can be disproportionately influenced by extreme values, meaning a model with a single large maximum error could misleadingly appear to have higher accuracy despite performing worse overall. By focusing on absolute errors, we provide a more stable and reliable comparison between the two models.
Across nearly all datasets and years, our spatial Kalman filter outperformed the generic Kalman filter, yielding lower absolute errors in every case except for the 2020 predictions of prescription dispensing rates. This exception is notable because the dispensing rates dataset exhibited substantial variability and lacked a clear spatial structure, perhaps making it less suited for a spatially informed modeling approach. The spatial Kalman filter is designed to capture underlying geographic correlations, and its advantage is most pronounced when such patterns exist. This suggests that while the spatial Kalman filter is a powerful tool for modeling spatiotemporal processes, its effectiveness may depend on the presence of meaningful spatial patterns in the data.
Figure 9 in appendix G presents a comparison of error distributions for four years in the study period, demonstrating the superior performance of our spatial Kalman filter in all but the case discussed above. Additional histogram comparisons for the full study period can be found in the accompanying GitHub repository. Overall, our spatial Kalman filter proves to be a more effective modeling approach than its generic counterpart. By integrating both temporal and geographic dependencies, the spatial Kalman filter enables more precise estimations, making it a valuable tool for studying complex datasets with inherent spatial patterns. The methodology we present can be broadly applied to other datasets where spatial relationships play a critical role and has broader utility beyond the opioid-related outcomes analyzed in this study.
Having established the spatial Kalman filter as a more effective framework than its generic counterpart, we now turn to the opioid vulnerability profiles that emerged from its predictions. The analysis of profiles for the mortality rates revealed widespread national vulnerability where the Appalachian region was repeatedly distinguished as the nation’s most vulnerable area to opioid-related mortality rates. Correspondingly, the analysis examining the disability rank ordered rate profiles also uncovered widespread vulnerability where Appalachia was again repeatedly pinpointed as the nation’s most vulnerable area to disability rank ordered rates. This alignment in the patterns for both disability and mortality profiles suggests that amidst the general fabric of the opioid crisis, the disabled subpopulation may be more at risk of opioid-related mortality than the general population. This finding corroborates previous work [33, 31, 55], and underscores the disproportionate outcomes of the opioid crisis.
@@ 5   Conclusion
In this study, we leveraged the predictions from our spatial Kalman filter to investigate opioid vulnerability profiles and identify critical spatiotemporal patterns. Our findings highlight the potential of this approach for capturing geographic dependencies in opioid-related outcomes and documenting shifts in vulnerability over time. While our model was specifically applied to opioid-related mortality, prescription dispensing, and disability rates, the methodologies developed here can be adapted for a broader range of applications involving spatial and temporal dynamics.
Our current framework models opioid-related mortality, prescription dispensing, and disability rates as independent processes, capturing temporal trends and spatial dependencies within each outcome separately. While incorporating socioeconomic covariates—such as poverty levels, healthcare access, or employment rates—could enhance interpretability, doing so within the current framework would model each outcome independently as a function of these factors, without explicitly capturing their interdependencies. A more comprehensive approach would involve a multivariate framework, in which socioeconomic factors serve as shared covariates across all three outcomes, enabling a more holistic analysis of their relationships. Future work could explore multivariate spatial Kalman filters, extended Kalman filters, or hierarchical modeling approaches to jointly model these interdependencies and better capture the complex interactions driving opioid vulnerability [58].
Finally, while our spatial Kalman filter effectively captures geographic proximity, temporal trends, and opioid-related outcomes, it does not explicitly adjust for external confounders such as policy interventions at the local, state, or federal levels. However, the model remains valuable for tracking shifts in opioid vulnerability, as external policy changes and other interventions will manifest in its predictions, revealing alterations in vulnerability patterns at multiple geographic scales—including the county, state, and national levels. It provides a critical tool for understanding how the opioid landscape evolves following such changes. Additionally, incorporating spatial correlations allows the model to capture regional patterns that traditional time-series methods like ARIMA or the generic Kalman filter may overlook, underscoring its broader applicability to public health research and spatial epidemiology.
Despite its strengths, this study does have certain limitations. The restricted range of years with consistent data across all three datasets posed challenges for capturing longer-term trends. The dispensing rates dataset contained a notable amount of missing data, requiring careful handling. Furthermore, the final year of this study’s time frame coincides with the onset of the COVID-19 pandemic, confounding the pandemic’s broader societal effects with the societal drivers of opioid abuse. These effects may limit the comparability of 2020 data to prior years and affect the generalizability of findings from that year.